9. . A ball rolling on a table has a velocity with rectangu- lar components vx 0.60 m/s and vy 0.80 m/s. What is the displacement of the ball in an interval of 2.5 s?
Answer #1

d = v t

dx = vx t = (0.60)(2.50) = 1.5 m

dy = vy t = 2 m

So displacement is 1.5 m along x axis and 2 m along y axis.

magnitude = sqrt(dx2 + dy2 ) = 2.5 m

angle with x aixs = tan^-1(2/1.5) =53 deg